OpenStudy (anonymous):

A dog chases a squirrel at a speed of 12 m/s. The squirrel dashes up a tree trunk at a rate of 6 m/s. Find the rate of change of distance between the squirrel and the dog at an instant when the do is 12m from the tree trunk and the squirrel is 5m up the trunk.

OpenStudy (anonymous):

ok. to set up this equation, use the pathagorean theorum. a^2+b^2=c^2 that gives you just about everything you need. other than that, take the derivitive. the key is understanding what each part means. rate of change is your key.

OpenStudy (anonymous):

Take the derivative of pythagorean's theorum?

OpenStudy (anonymous):

basically yes. taking the derivitive would be 2ada+2bdb=2cdc given a=12, b=5, you can find c from the original equation 144+25=c^2, c=13 da=12, db=6, so 2*12*12+2*5*6=2*13*dc 288+60=26dc 348=26dc 13.385=dc